OpenStudy (anonymous):

Write the product as a trinomial. (x + 2)(x + 4)

OpenStudy (anonymous):

(x+2)(x+4) =(x+2)(x+4) =(x)(x)+(x)(4)+(2)(x)+(2)(4) =x2+4x+2x+8 =x2+6x+8

OpenStudy (anonymous):

it is better to write x^2 +6x +8 x^2 means to the power ie 2^2 =4 the upside down v means power the other way can be confused with multiply